Quect Shipping Company charges a flat fee of $6 plus an additional $0.14 per pound to ship a package. Coraz Shipping Company charges a flat fee of $12 plus an additional $0.10 per pound. How many pounds would a package weigh that would cost the same at both companies.
1
Expert's answer
2020-12-15T02:12:45-0500

For first company cost is:


"c_1=6+0.14x",


where "x" - weight of a package in pounds.


For second company cost is:


"c_2=12+0.1x".


If costs are the same, then:


"6+0.14x=12+0.1x;\\
0.04x=6;\\
x=150."


Answer: 150 pounds